OpenStudy (anonymous):

Need help with Taylor expansion of an integral: integral from o to 1 of Sin(t)/t I am having trouble setting this Taylor series up...for example, will the a=1? I know I need to use (x-a)^n/n! to set up the polynomials, but I'm not sure what a would be.

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\sin t/t \]

OpenStudy (turingtest):

nope, I don't think I know what we're talking about here...

OpenStudy (turingtest):

oh but apparently I just went up a level for typing that remark lol

OpenStudy (anonymous):

Oh also, when solving for the coefficients, should I plug in zero or 1?

hero (hero):

Welcome to level 99

OpenStudy (turingtest):

haha

OpenStudy (anonymous):

@mdntjem u just need to write taylor expansion of \(\sin t\) for \(-\infty<t<\infty\) \[\sin t=\sum_{n=0}^{\infty} (-1)^n\ \frac{t^{2n+1}}{(2n+1)!}\]

OpenStudy (turingtest):

why does it talk about the integral then?

OpenStudy (turingtest):

I don't see the relation...

OpenStudy (anonymous):

so when I find the dervatives, I just use sinx not sinx/x?

OpenStudy (anonymous):

well... i think she wants to represent the integral \[\int \frac{\sin t}{t}dt\] in the form of a series

OpenStudy (anonymous):

Yep, that's what the integral is supposed to look like, but it's missing the limits.

OpenStudy (turingtest):

so you still want to integrate the series term by term or what?

OpenStudy (turingtest):

because I suppose you could do that, right?

OpenStudy (anonymous):

@mdntjem divide by t that expression mukushla mentioned above.

OpenStudy (turingtest):

...and then integrate it, yes?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

ok that makes some sense but how do you evaluate it as a definite integral?

OpenStudy (anonymous):

answer will be a series again...

OpenStudy (anonymous):

Oh I did not know you could do that, thanks! Yea, I just need to find the first 3 NON-zero terms of the series and the add em up

OpenStudy (anonymous):

its easy . it depends on the margin of error. she can integrate a few terms. (it will be series but she can add the terms to get value ). there will be some error associated.

OpenStudy (anonymous):

but I am still confused would my a=1 when evaluating the polynomials? And would this be centered around zero since I was not give a value?

OpenStudy (anonymous):

I think the question should have been to find the taylor approximation about a=0 (maclaurin series).otherwise it would not be possible to approximate the given integral . Can you check your Question in the book again.?

OpenStudy (anonymous):

i think she has problem with choosing \(a\) for writing taylor expansion...she can use both \(a=0\) and \(a=1\) (or any other neighborhood) for evaluating the integral.... better to use \(a=0\) brcause obtaining the coefficients will be very easy about \(a=0\)...

OpenStudy (anonymous):

Okay, your guys are right, it does not make sense for a=1 I just wrote it all out

OpenStudy (anonymous):

Because I need to plug 1 in for x in order to evaluate the series

OpenStudy (anonymous):

Thank you all for your help!